# SLIFER: Investigating Performance and Robustness of Malware Detection Pipelines

### v1

@@ 5.4. Computation Times Comparison
We now analyse the classification times of all the modules and models tested, and we report our findings in Table 7 to answer RQ4. We measure MDT for malware and benignware separately on both $\mathcal{D}_{1}$, and $\mathcal{D}_{2}$, as explained in Sect. 5, including the standard deviation of each measurement as well. In general, we can see a difference between times computed for malware (first column) and benignware (second column). This discrepancy is mainly attributed to the larger dimension of the benignware in terms of bytes and content included in programs, thus slowing down all the compared techniques. As expected, signatures and static models are the fastest techniques, with MalConv being the best one, followed by YARA pattern-matching and GBDT-EMBER. Intuitively, MalConv does not pre-process input samples, and it just uses their first 1MB to compute predictions without relying on feature extractors. On the contrary, while pattern-matching is indeed fast, it is slowed down by the length of the file to analyse, requiring different amounts of computations for smaller and larger programs. Lastly, GBDT-EMBER is two orders of magnitude slower than MalConv due to the heavy feature extraction phase encoded through EMBER. Unsurprisingly, hybrid and dynamic analysis with Quo.Vadis and Nebula requires a non-negligible amount of seconds due to emulation. On the contrary, even if built on top of many components, SLIFER rivals GBDT-EMBER when analysing malware. Since computations are halted at first detection, SLIFER cuts the heavy feature-extraction phase when possible. On the contrary, since benign samples must traverse all the pipeline, we report a slightly higher required time to analyse benignware, mostly due to the emulation step. Thus, we can conclude that the overhead induced by SLIFER is negligible w.r.t. both static and dynamic analysis alone.
@@
Table 7. Mean Detection Time (MDT) of each module or architecture testes, separately computed for malware and benignware subsets of the original datasets.
@@
Take-home message 4: Classification times strongly depend on the size of files, and the presence of emulation. The architecture of SLIFER does not introduce noticeable overheads while keeping all types of analyses in one place.
6. Related Work
@@
Hybrid analysis. In recent years, academia has focused on many architectures for ML malware detectors that mix both static and dynamic analysis (Gibert et al., 2020). After the establishment of Machine Learning models in the malware analysis domain, research focused on how to train models extracting features from these two analysis approaches to increase performance in binary classification and/or family classification. As already mentioned in this work, with Quo.Vadis (Trizna, 2022), Trizna developed a multiple deep learning models architecture (early fusion models) that analyse both static and dynamic features and then merges results with a single meta-model. This approach is well-known in literature as late fusion strategy. Han et al. (Han et al., 2019a, b) tested different ML models in malware detection and classification tasks, merging static and dynamic features, and training models directly with the whole set of mixed features. The same approach is described by Kumar et al. (Kumar et al., 2019). All these works apply the so-called early fusion strategy, and experiments report an improvement in accuracy. Ngo et al. (Ngo et al., 2023) tackled the problem of feature fusion with Transfer Learning and Knowledge Distillation (KD): a large teacher model was trained on aggregated static and dynamic features and a small student model only on static features. Then the knowledge of the rich ”behavior-aware” model is transferred to the faster small model to perform classification. Similarly to our work, this approach aims at reducing detection delays caused by dynamic analysis: indeed, after the training effort of the teacher model including behavioral features, this knowledge is only transferred to the student model, which computes only static features to perform classification.
@@
There are differences between related works and ours: differently from the techniques mentioned before, we also focus on evaluating the robustness of SLIFER to adversarial EXEmples, analysing the contribution of dynamic model and signatures on ADR; moreover, we present two different policies for impossible-to-analyse samples, while in literature they are discarded or they are not the object of interest for any further analysis.
7. Limitations
We now evaluate the limitations of our work, by also discussing how they can be handled or why they are not relevant.
@@ 7. Limitations
SLIFER Model Training. Our pipeline’s architecture is composed of pre-trained objects on different datasets. The scope of our work is not to build a new model to optimize a common loss function, training the whole pipeline on the same dataset. Still, we want to highlight the performances of a sequence of ML models in detecting malware, adding signatures to static and dynamic analysis, similarly to industrial architectures, analysing the contributions of each module. Building a single model which can analyse different features without relying on feature fusion (as Quo.Vadis (Trizna, 2022)) is a non-trivial problem out of the scope of this work.
@@
Dynamic Analysis on Virtualization. Nebula model is trained on reports generated through Speakeasy (Mandiant. Speakeasy: portable, modular, binary emulator designed to emulate Windows kernel and user mode malware, 2022), that is an emulation tool. We see limitations in PE emulation, especially in reliability, as assessed by our experimental work. Virtualizing a huge amount of PE as done for the Speakeasy dataset (Trizna, 2022) needs time and a well-settled environment and it is out of our purpose, which is testing available ML models for dynamic analysis inside a sequential pipeline.
@@
End-to-End Adaptive Attacks. We perform transfer attacks targeting only GBDT, namely the best-performing one. Attacks on ML models trained on static features are well implemented and documented (Demetrio et al., 2021a). Evading dynamic analysis is a wide problem (Afianian et al., 2019) and currently, we lack an implementation of attacks to evade dynamic models. As far as we know, there is only one attack that has been proposed (Rosenberg et al., 2018), but no code has been provided. Moreover, we do not customise GAMMA to avoid including any artifact detected by YARA rules. However, our goal for this paper is the empirical analysis of malware pipelines that resembles industry settings, without delving into adversarial robustness, as this would require a deeper investigation on its own, as well as developing novel or re-implementing attacks.
8. Conclusions
@@
In this work, we propose SLIFER, a novel Windows malware detector pipeline, that sequentially puts together static and dynamic analysis, leveraging pattern-matching with YARA and state-of-the-art machine learning models for malware detection. We build an architecture that mimics industrial solutions, testing its performances compared to hybrid approaches, proposing policies to handle analysis errors, and analysing its robustness against adversarial EXEmples. We highlight interesting findings remarked by our experimental work: SLIFER proves to outperform state-of-the-art models designed for static, dynamic, or hybrid analysis, keeping the highest TPR with low FPR. Moreover, it proves faster detection times of malicious PEs as compared to models leveraging dynamic analysis; in achieving this, our proposed architecture does not discard any input sample and propagates detection errors until the end of the pipeline, and labeling impossible-to-analyse samples as benign proves to be the best policy to provide less false alarms. SLIFER shows its robustness against adversarial transfer attacks and surprisingly demonstrates that section injection attacks are detected more by YARA rules than by dynamic analysis, whose integrity is mined by small changes in the sample structure that are reflected in the emulation report. Overall, we report the small contribution of dynamic analysis based on emulation to accuracy and robustness, besides the computational burden that is required, proving the effectiveness of our choice to use that kind of analysis only when required and in the last resort.
@@
Future Work. As mentioned in Sect. 7, we did not learn all the parameters of the components of SLIFER, but rather we leveraged pre-trained models. However, we believe that it is possible to formalize an algorithm that wraps all modules, whose parameters can be learned within a single minimization by also keeping static and dynamic analyses separated. Also, while in Another future investigation will take into account the difference between emulation and virtualization. In particular, we will train the dynamic analysis module on virtualization reports collected from a training PE dataset, and we willniivestigate the difference in both performance and overhead. As regards end-to-end attacks, we plan on developing novel adversarial attacks suited to evade end-to-end malware pipelines like SLIFER, thus trying to deceive both malware signatures and dynamic modules.
Acknowledgements.
@@
Andrea Ponte acknowledges the support of Rina Consulting Spa to his doctoral scholarship and research work. This work was partially supported by project SERICS (PE00000014) under the NRRP MUR program funded by the EU - NGEU. This work was partially supported by project FAIR (PE00000013) under the NRRP MUR program funded by the EU - NGEU.

### v3

@@ 5.5 Computation Times Comparison
We now analyze the classification times of all the modules and models tested, also considering the calibrated SLIFER, and we report our findings in Table 12 to answer RQ6. We measure MDT for malware and benignware separately on both $\mathcal{D}_{1}\,$, and $\mathcal{D}_{2}\,$, as explained in Sect. 5, including the standard deviation of each measurement as well. In general, we can see a difference between times computed for malware (first column) and benignware (second column). This discrepancy is mainly attributed to the larger dimension of the benignware in terms of bytes and content included in programs, thus slowing down all the compared techniques. As expected, signatures and static models are the fastest techniques, with MalConv being the best one, followed by YARA pattern-matching and GBDT-EMBER. Intuitively, MalConv does not pre-process input samples, and it just uses their first 1MB to compute predictions without relying on feature extractors. On the contrary, while pattern-matching is indeed fast, it is slowed down by the length of the file to analyze, requiring different amounts of computations for smaller and larger programs. Lastly, GBDT-EMBER is two orders of magnitude slower than MalConv due to the heavy EMBER feature extraction phase. Unsurprisingly, hybrid and dynamic analysis with Quo.Vadis and Nebula requires a non-negligible amount of seconds due to emulation. On the contrary, even if built on top of many components, SLIFER rivals GBDT-EMBER when analysing malware. Since computations are halted at first detection, SLIFER cuts the heavy feature-extraction phase when possible. On the contrary, since benign samples must traverse all the pipeline, we report a slightly higher required time to analyze benignware, mostly due to the emulation step. Thus, we can conclude that the overhead induced by SLIFER is negligible w.r.t. both static and dynamic analysis alone. However, on calibrated SLIFER, we highlight that, on average, samples require more time to be analyzed on average. Since MalConv decision threshold increases, more malicious samples are sent to the next modules. At the same time, due to the reduction of false positives, more benignware do not trigger false alarms and they are analyzed by all the components of SLIFER. Together, these two facts lead to an increase in the average analysis time, both for benign and malicious samples.
Table 12: Mean Detection Time (MDT) of each module or architecture testes, separately computed for malware and benignware subsets of the original datasets.
@@
After Calibration
SLIFER No Dyn. $1.2\pm 3.4\times 10^{-1}$ $3.9\pm 9.7\times 10^{-1}$
SLIFER $2.3\pm 14\times 10^{-1}$ $6.6\pm 19$
@@
Take-home message 6: Classification times strongly depend on the size of files, and the presence of emulation. The architecture of SLIFER does not introduce noticeable overheads while keeping all types of analyzes in one place. However, calibration induces a slightly-higher analysis time, due to the change of which module detect which malware.
6 Related Work
After the establishment of Machine Learning models in the malware analysis domain, academic research has concentrated on analyzing ML models for malware detection, introducing numerous architectures for Machine Learning-based malware detectors that utilize signatures, static features, dynamic features, or a combination of them, known as hybrid analysis Gibert et al. (2020); Ye et al. (2017). Our work can be labeled as the latter, which complements the information retrieved from static analysis with behavioral aspects retrieved during dynamic analysis. While our work focuses on sequential analysis, which keeps each analysis separate by submitting samples to different classifiers (each trained on distinct features) in a step-by-step sequence, previous work considered mixing both static and dynamic features mainly in two ways. Early fusion strategy merges all the features before training a single classifier. Instead, in late fusion strategy, separate classifiers for each family of features are trained and then the results are combined in a single score. We summarize in Table 13 a pool of the most relevant works that studied how to combine features from different malware analyzes, by annotating each work with the applied hybrid analysis employed (early fusion, late fusion or sequential), the source of the selected features (from malware signatures, static or dynamic analysis) and we report the analysis presented to assess the quality of the described architectures. For this last column, we report keywords for brevity and readability.
@@ 6 Related Work
As already mentioned in this work, with Quo.Vadis, Trizna (2022) developed a deep learning model architecture that analyze separately static and dynamic features and then merges results with a single meta-model through late fusion. Yen et al. (2019) mix several features computed from both static and dynamic analysis to build early fusion models, late fusion models, and then a mixed fusion neural network. Also Ma et al. (2016) use an ensemble of different classifiers to perform a hybrid analysis building training models with separate feature sets made coupling static and dynamic analysis telemetry. Shijo and Salim (2015) document a way to construct a feature vector leveraging data from static and dynamic analysis techniques and processed by a single ML model.  Han et al. (2019a) and  Han et al. (2019b) tested different ML models in malware detection and classification tasks, merging static and dynamic features, and training models directly with the whole set of mixed features. The same approach is described by Kumar et al. (2019), which also report an interesting analysis on virtualization times: they study how accuracy varies adjusting the run time of samples in a sandbox. All these works apply the so-called early fusion strategy, and experiments report an improvement in accuracy using both families of features.  Ngo et al. (2023) tackled the problem of feature fusion with Transfer Learning and Knowledge Distillation (KD): a large teacher model was trained on aggregated static and dynamic features and a small student model only on static features. Then the knowledge of the rich "behavior-aware" model is transferred to the faster small model to perform classification. Similarly to our work, this approach aims at reducing detection delays caused by dynamic analysis: indeed, after the training effort of the teacher model including behavioral features, this knowledge is only transferred to the student model, which computes only static features to perform classification.
@@
Different from most of the works mentioned before, we also focus on evaluating the robustness of SLIFER to adversarial EXEmples, analyzing the contribution of dynamic model and signatures on ADR; moreover, we present two different policies for impossible-to-analyze samples, while in literature they are discarded or they are not the object of interest for any further analysis. Moreover, as highlighted by Table 13, as far as we know, we are the first to analyze a sequential technique that chains together different kinds of analysis, considering error-handling, robustness, and calibration of those modules.
Table 13: Overview of Related Works described in Sect. 6. We describe the type of hybrid analysis methodology employed, the sources of features and malware characteristics used to construct the dataset. We also provide the keywords for the analyzes performed in the experimental evaluation to test the proposed architectures.
Methodology Malware Analysis Type Discussed Analysis
Work Early Fusion Late Fusion Sequential Signatures Static Dynamic Keywords
Trizna (2022) Performance, Robustness
Yen et al. (2019) Performance
Shijo and Salim (2015) Performance
Ma et al. (2016) Performance
Han et al. (2019a) Performance
Han et al. (2019b) Performance
Kumar et al. (2019) Performance, Detection Times, Virtualization Times
Ngo et al. (2023) Performance, Detection Times
SLIFER Performance, Error Handling, Robustness, Model Calibration, Detection Times
7 Limitations
We now evaluate the limitations of our work, by discussing whether to handle them or their irrelevance.
@@ 7 Limitations
SLIFER Model Training. Our pipeline’s architecture is composed of pre-trained objects on different datasets. The scope of our work is not to build a new model to optimize a common loss function, training the whole pipeline on the same dataset. Still, we want to highlight the performances of a sequence of ML models in detecting malware, adding signatures to static and dynamic analysis, similarly to industrial architectures, analysing the contributions of each module. Building a single model which can analyze different features without relying on feature fusion as Quo.Vadis (Trizna, 2022) is a non-trivial problem out of the scope of this work.
SLIFER Calibration. In this work, we leverage pre-trained models and we propose a tentative fine-calibration of decision thresholds. In doing so, we perform an exhaustive grid search with different values models’ thresholds inside SLIFER. As mentioned above, we do not have a common loss function to minimize in an end-to-end learning process, and this does not let us perform a proper cross-validation or other fine-tuning techniques.
@@
Evaluation of Analysis Errors. In our analysis, we report that labeling errors as benign helps reduce FPR, as demonstrated by experimental work. However, in a real-world scenario this could be harmful since an error can bypass SLIFER’s detection in its actual configuration. While in this work we propose a first basic policy of error handling, we discuss more fine-grained solutions in Sect. 8.
@@
Dynamic Analysis on Virtualization. Nebula model is trained on reports generated through Speakeasy (Mandiant. Speakeasy: portable, modular, binary emulator designed to emulate Windows kernel and user mode malware, ), that is an emulation tool. We see limitations in PE emulation, especially in reliability, as assessed by our experimental work. Virtualizing a huge amount of PE as done for the Speakeasy dataset (Trizna, 2022) needs time and a well-settled environment and it is out of our purpose, which is testing available ML models for dynamic analysis inside a sequential pipeline.
@@
End-to-End Adaptive Attacks. We perform transfer attacks targeting only GBDT, namely the best-performing one. Attacks on ML models trained on static features are well implemented and documented (Demetrio et al., 2021a). However, while the evasion of dynamic analysis is a wide and well-known problem (Afianian et al., 2019), currently no open source code has been released to re-evaluate previously-proposed techniques. As far as we know, there is only one attack that has been proposed by (Rosenberg et al., 2018), but it is not replicable since no code has been provided. Moreover, we do not customise GAMMA to avoid including any artifact detected by YARA rules. However, our goal for this paper is the empirical analysis of malware pipelines that resembles industry settings, without delving into adversarial robustness, as this would require a deeper investigation on its own, as well as developing novel or re-implementing attacks.
8 Conclusions
@@
In this work, we are the first to study the behavior of sequential malware detection pipelines in terms of performance, resilience to errors, and robustness. To do so, we develop SLIFER, a Windows malware detector pipeline, that sequentially puts together static and dynamic analysis, leveraging pattern-matching with YARA and state-of-the-art machine learning models for malware detection. We build an architecture that mimics industrial solutions, testing its performances compared to hybrid approaches, proposing policies to handle analysis errors, and analysing its robustness against adversarial EXEmples. We highlight interesting findings remarked by our experimental work: SLIFER proves to outperform state-of-the-art models designed for static, dynamic, or hybrid analysis, keeping the highest TPR with low FPR. Moreover, it proves faster detection times of malicious PEs as compared to models leveraging dynamic analysis; in achieving this, our proposed architecture does not discard any input sample and propagates detection errors until the end of the pipeline, and labeling impossible-to-analyze samples as benign proves to be the best policy to provide less false alarms. Moreover, we present a simple way to calibrate decision thresholds of a series of pre-trained models, performing an exhaustive grid search on a validation set. We state that the calibration effort can bring advantages in terms of TPR and FPR in most scenarios, minimally sacrificing fast detection times. SLIFER exhibits robustness against adversarial transfer attacks, also surprisingly demonstrates that GAMMA section-injection is detected more by YARA rules than by dynamic analysis, whose integrity is mined by small changes in the sample structure that are reflected in the emulation report. Overall, we report the small contribution of emulation-based dynamic analysis to accuracy and robustness, besides the computational burden that is required. This proves of the effectiveness of our choice to use that kind of analysis only when required and as the last resort. Lastly, we underline the benefits of calibration on robustness, highlighting how SLIFER blocks more adversarial EXEmples generated with section injection and padding thanks to the tuning of the thresholds of the pre-built GBDT-EMBER, MalConv, and Nebula.
@@
Future Work. As mentioned in Sect. 7, we did not learn all the parameters of the components of SLIFER, but rather we leveraged pre-trained models. However, we believe that it is possible to formalize an algorithm that wraps all modules, whose parameters can be learned within a single minimization by also keeping static and dynamic analyzes separated. In addition, we want to improve the calibration mechanism, including other parameters in the grid, or other metrics to compute on the validation set, such as the detection time. Or, after finding a process to perform end-to-end training, we can perform proper cross-validation or other fine-tuning techniques. We would want to improve the empirical error handling technique we deployed, by considering two alternatives. The first one considers relying on end-to-end models like MalConv as they are error-proof due to the lacking of feature extraction process. Hence, we could design an ensemble of end-to-end learners to give a more accurate decision on impossible-to-analyze samples. Thus, if further components of the pipeline would get an error, the decision would be taken by this ensamble instead. Connected to this solution, a second alternative is to implement conformal prediction (Vovk et al., 2005; Shafer and Vovk, 2008). Error-proof models can be tuned using a conformal prediction framework (Angelopoulos et al., 2023) to know when (at which probability threshold) such models are certain in labeling a sample as benign, thus reducing the numbers of later errors. Also, another future investigation will take into account the difference between emulation and virtualization. In particular, we will train the dynamic analysis module on virtualization reports collected from a training PE dataset, and we will investigate the difference in both performance and overhead. As regards end-to-end attacks, we plan on developing novel adversarial attacks suited to evade end-to-end malware pipelines like SLIFER, thus trying to deceive both malware signatures and dynamic modules.
Acknowledgments
@@
Andrea Ponte acknowledges the support of Rina Consulting S.p.A. to his doctoral scholarship and research work. This work was partially supported by projects SERICS (PE00000014) and FAIR (PE00000013) under the NRRP MUR program funded by the EU - NGEU.